# MoH: Multi-Head Attention as Mixture-of-Head Attention

### v2

@@ Abstract
In this work, we upgrade the multi-head attention mechanism, the core of the Transformer model, to reduce computational costs while maintaining or surpassing the previous accuracy level. We show that multi-head attention can be expressed in the summation form. Drawing on the insight that not all attention heads hold equal significance, we propose Mixture-of-Head attention (MoH), a new architecture that treats attention heads as experts in the Mixture-of-Experts (MoE) mechanism. MoH has two significant advantages: First, MoH enables each token to select the appropriate attention heads, enhancing inference efficiency without compromising accuracy or increasing the number of parameters. Second, MoH replaces the standard summation in multi-head attention with a weighted summation, introducing flexibility to the attention mechanism and unlocking extra performance potential. Extensive experiments on ViT, DiT, and LLMs demonstrate that MoH outperforms multi-head attention by using only 50$\sim$90of the attention heads. Moreover, we demonstrate that pre-trained multi-head attention models, such as LLaMA3-8B, can be further continue-tuned into our MoH models. Notably, MoH-LLaMA3-8B achieves an average accuracy of 64.0% across 14 benchmarks, outperforming LLaMA3-8B by 2.4% by utilizing only 75% of the attention heads. We believe the proposed MoH is a promising alternative to multi-head attention and provides a strong foundation for developing advanced and efficient attention-based models. The code is available at https://github.com/SkyworkAI/MoH.
@@ 1 Introduction
Since attention is introduced and becomes a fundamental component of Transformers (Vaswani et al., 2017), multi-head attention has been the standard architecture for natural language processing (Kenton & Toutanova, 2019) and computer vision tasks (Dosovitskiy et al., 2021). It is well known that using multiple heads can improve model accuracy. However, not all attention heads hold equal significance. Some works have shown that many attention heads can be pruned without affecting accuracy. For example, Voita et al. (2019) introduces a method to quantify the usefulness of each attention head and prune those that are redundant. Similarly, Michel et al. (2019) challenges the necessity of multiple heads by examining the impact of extensive pruning across various settings. In computer vision, some works also identify attention head redundancy. Bhattacharyya et al. (2023) reduces redundancy to boost performance, while Yun & Ro (2024) develop single-head attention for efficiency. These findings demonstrate that vanilla multi-head attention contains redundant attention heads.
@@
To this end, we introduce Mixture-of-Head attention (MoH), a new architecture that integrates multi-head attention with the Mixture-of-Experts (MoE) mechanism (Jacobs et al., 1991). Specifically, we propose to treat attention heads as experts within the MoE framework. Similar to MoE, MoH consists of multiple attention heads and a router that activates the Top-K heads for each token. Moreover, we replace the standard summation in multi-head attention with a weighted summation. This design offers two significant advantages: First, MoH allows each token to select the most relevant attention heads, improving inference efficiency without sacrificing accuracy or increasing the parameters. Second, by replacing the standard summation in multi-head attention with a weighted summation, MoH enhances the flexibility of the attention mechanism and increases the performance potential. Moreover, to efficiently capture common knowledge across different contexts, we designate a subset of attention heads as shared heads that remain always activated.
We evaluate our proposed MoH across various popular model frameworks, including Vision Transformers (ViT) (Dosovitskiy et al., 2021) for image classification, Diffusion models with Transformers (DiT) (Peebles & Xie, 2023) for class-conditional image generation, and Large Language Models (LLMs) (Brown et al., 2020; OpenAI, 2022; Ouyang et al., 2022). We show that MoH achieves competitive performance, or even outperforms multi-head attention with only 50$\sim$90of the attention heads. For example, MoH-ViT-B achieves 84.9%/84.7% Top-1 accuracy on the ImageNet-1K (Deng et al., 2009) classification benchmark, surpassing well-tuned multi-head attention baselines with only 75%/50% of the attention heads.
Furthermore, we demonstrate that pre-trained multi-head attention models, such as LLaMA3-8B (Dubey et al., 2024), can be further continue-tuned into our MoH models. Specifically, using only about 3% (400B tokens) of the original LLaMA3 pre-training data for continue-tuning, MoH-LLaMA3-8B achieves an average accuracy of 64.0% across 14 benchmarks, outperforming LLaMA3-8B by 2.4% by utilizing only 75% of the attention heads. These results show that MoH is a promising alternative to vanilla multi-head attention, laying a solid foundation for developing advanced and efficient attention-based models. The main contributions are summarized as follows:
@@ 2 Related Work
Multi-Head Attention.    Transformers (Vaswani et al., 2017) have garnered significant interest and success in both natural language processing and computer vision. The success of transformers has been long attributed to the multi-head attention mechanism (Cordonnier et al., 2020). Multi-head attention mechanism is proposed by Vaswani et al. (2017) to enhance the representation power of an attention layer by allowing multiple attention heads to operate on different low-dimensional projections of the input. The outputs from these heads are then concatenated to form the final result. Alternatively, by decomposing the output projection matrix by rows, multi-head attention can be expressed in a summation form. In summation form, each head operates in parallel, and the final output is the sum of all heads. Inspired by this observation, we propose MoH, a dynamic attention-head routing mechanism that allows each token to adaptively select the appropriate heads.
@@
Mixture-of-Experts Models.    The Mixture-of-Experts (MoE) method (Du et al., 2022; Lewis et al., 2021; Rajbhandari et al., 2022; Roller et al., 2021; Zhou et al., 2022; Jin et al., 2025) is introduced to expand the capacity of deep neural networks without increasing computational costs. In this approach, only a subset of parameters, known as experts, is activated for each input. Shazeer et al. (2017) first introduces an MoE layer between LSTM layers. Switch Transformer (Fedus et al., 2022) further simplifies the gating mechanism by selecting only the Top-1 expert per token. Gshard (Lepikhin et al., 2021) improves the Top-2 expert routing strategy. In contrast to MoE, which emphasizes efficient parameter scaling while maintaining manageable computational costs, our MoH focuses on reducing the activation of redundant attention heads without increasing the number of parameters.
@@
Attention Head Specialization and Efficiency.    Many recent studies show that not all attention heads in Transformers are equally useful. Peng et al. (2020) proposes a mixture-of-heads approach, where only a few selected heads are used, yet the model performs just as well or even better. Csordás et al. (2024) pushes this further with SwitchHead, an MoE-style method that activates only a small number of heads for each token, speeding up inference while keeping performance high. In long-context language models, this idea is even more clear. Wu et al. (2024) shows that a few special retrieval heads are mainly responsible for keeping facts consistent in long inputs. Fu et al. (2024) finds that keeping only the most useful heads in the KV cache can save memory. Xiao et al. (2024) proposes DuoAttention, which combines different types of heads to make long-context inference more efficient without losing quality. Similar patterns appear in vision models. Gandelsman et al. (2023) shows that CLIP’s attention heads each focus on specific visual features, and this can be explained through related text prompts. Balasubramanian et al. (2024) finds that this kind of head specialization also exists in other vision models beyond CLIP. Basile et al. (2024) shows that using only a few selected heads chosen by spectral methods can even beat the full model on zero-shot tasks.
@@ 3.1 Multi-Head Attention
We begin by reviewing the multi-head attention mechanism introduced by Vaswani et al. (2017). The multi-head attention mechanism is based on scaled dot-product attention. Specifically, for $T$ tokens $\bm{X}\in\mathbb{R}^{T\times d_{in}}$ of $d_{in}$ dimensions each and $T^{\prime}$ tokens $\bm{X}^{\prime}\in\mathbb{R}^{T^{\prime}\times d_{in}}$ of $d_{in}$ dimensions each, the scaled dot-product attention is computed as follows:
@@
Concatenation Form.    To enhance the representation power, Vaswani et al. (2017) proposes to allow multiple attention heads to operate on different low-dimensional projections of the input tokens. Specifically, the multi-head attention mechanism computes $h$ different low-dimensional projections of $(\bm{Q},\bm{K},\bm{V})$, performs scaled dot-product attention for each head, concatenates the results, and applies a projection to the concatenated output. The concatenation form of the multi-head attention can be formulated as:
[…]
@@
Summation Form.    The multi-head attention mechanism is typically represented in its concatenation form. However, from another perspective, if we decompose $\bm{W}_{O}\in\mathbb{R}^{d_{v}\times d_{out}}$ by rows, we can express multi-head attention in a summation form. Specifically, $\bm{W}_{O}$ can be divided into $h$ matrices by rows, i.e., $[\bm{W}_{O}^{1},\bm{W}_{O}^{2},...,\bm{W}_{O}^{h}]=\bm{W}_{O}$, where $\bm{W}_{O}^{i}\in\mathbb{R}^{d_{v}/h\times d_{out}}$. Finally, the summation form of the multi-head attention can then be formulated as:
@@
Table 1: Comparisons to current state-of-the-art methods on ImageNet-1K classification. Our MoH-ViT models, based on TransNeXt (Shi, 2024), are trained for 300 epochs using a resolution of 224$\times$224. To ensure a fair comparison, we only replace the standard multi-head attention with our Mixture-of-Head attention (MoH), keeping all other training parameters identical to TransNeXt.
@@
DeiT-S (Touvron et al., 2021) 22 100 79.8
T2T-ViT-19 (Yuan et al., 2021) 39 100 81.9
Swin-S (Liu et al., 2021) 50 100 83.1
PVTv2-B3 (Wang et al., 2022) 45 100 83.2
CoAtNet-1 (Dai et al., 2021) 42 100 83.3
Focal-S (Yang et al., 2021) 51 100 83.5
FocalNet-S (Yang et al., 2022b) 50 100 83.5
MViTv2-S (Li et al., 2022) 35 100 83.6
UniFormer-B (Li et al., 2023b) 50 100 83.9
CAFormer-S36 (Yu et al., 2023) 39 100 84.5
TransNeXt-S (Shi, 2024) 50 100 84.7
@@
DeiT-B (Touvron et al., 2021) 86 100 81.8
T2T-ViT-24 (Yuan et al., 2021) 64 100 82.3
Swin-B (Liu et al., 2021) 88 100 83.5
PVTv2-B5 (Wang et al., 2022) 82 100 83.8
Focal-B (Yang et al., 2021) 90 100 83.8
FocalNet-B (Yang et al., 2022b) 89 100 83.9
CoAtNet-2 (Dai et al., 2021) 75 100 84.1
MViTv2-B (Li et al., 2022) 52 100 84.4
MOAT-2 (Yang et al., 2022a) 73 100 84.7
iFormer-L (Si et al., 2022) 87 100 84.8
TransNeXt-B (Shi, 2024) 90 100 84.8
@@ 3.2 Mixture-of-Head Attention
Recently, the Mixture-of-Experts (MoE) method has emerged as a popular approach for scaling the parameters of large language models (Jiang et al., 2024; Muennighoff et al., 2024). A MoE layer consists of multiple expert networks and a router that activates the Top-K experts. Generally, the number of activated experts $K$ is significantly smaller than the total number of experts to ensure inference efficiency.
@@
Shared Heads.    In attention mechanism, some attention heads may capture common knowledge across different contexts, such as grammatical rules in language. Inspired by Dai et al. (2024), we designate a subset of heads as shared heads that remain always activated. By consolidating common knowledge within shared heads, we reduce redundancy among the other dynamically routed heads.
@@
$[\alpha_{1},\alpha_{2}]=\text{Softmax}(\bm{W}_{h}\bm{x}_{t}),$ (6)
@@
Table 2: Comparisons to current state-of-the-art methods on the benchmarking of class-conditional image generation on ImageNet-1K at 256$\times$256 resolution.$\uparrow$” denotes that higher is better. $\downarrow$” denotes that lower is better. “cfg” denotes the classifier-free diffusion guidance scale. “400K” denotes the training budget is 400K training steps.
@@
Load Balance Loss    Directly training an MoE layer often causes the majority of tokens to be routed to a small number of experts, leaving the remaining experts insufficiently trained (Shazeer et al., 2017). To avoid the unbalanced load in the proposed MoH, following previous MoE methods (Lepikhin et al., 2021; Wei et al., 2024), we apply a load balance loss. Specifically, for the $t_{th}$ input token $\bm{x}_{t}\in\mathbb{R}^{d_{in}}$ in $\bm{X}\in\mathbb{R}^{T\times d_{in}}$, the load balance loss $\mathcal{L}_{b}$ is formulated as:
@@ 4.1 ViT for Image Classification
Model Settings.    For Vision Transformers (ViT) (Dosovitskiy et al., 2021), our MoH-ViT models are implemented based on the TransNeXt (Shi, 2024) framework and trained from scratch on the ImageNet-1K dataset (Deng et al., 2009), which contains over 1.2 million images in 1,000 categories. To ensure a fair comparison, we only replace the standard multi-head attention with the proposed MoH, while keeping all other training parameters identical to TransNeXt.
Training Details.    Our MoH-ViT models are trained for 300 epochs using automatic mixed precision across 8 GPUs. We follow the training strategy of TransNeXt, which includes various data augmentation techniques, including Random Augmentation (Cubuk et al., 2020), Mixup (Zhang, 2017), CutMix (Yun et al., 2019), and Random Erasing (Zhong et al., 2020). We also apply Label Smoothing (Szegedy et al., 2016) and DropPath (Huang et al., 2016) to regularize our models. We optimize our models using AdamW optimizer (Loshchilov & Hutter, 2017) with a gradient clipping norm of 1.0 and a weight decay of 0.05. The initial learning rate is set to 1e-3, with a 5-epoch warm-up starting at 1e-6. A cosine learning rate scheduler (Loshchilov & Hutter, 2016) is employed to decay the learning rate. During training, images are randomly cropped to a size of 224$\times$224. It is worth noting that we do not use Exponential Moving Average (EMA) weights.
Table 3: Comparisons between MoH-LLMs and vanilla LLMs.100B” denotes a training budget of 100 billion tokens, while “200B” denotes a budget of 200 billion tokens. We observe that larger models, e.g., MoH-LLM-B, generally perform worse than smaller models, e.g., MoH-LLM-S, on TruthfulQA, consistent with the findings reported by Lin et al. (2022).
@@
LLaMA3-8B (Dubey et al., 2024) 100 81.9 30.0 83.9 75.5 94.0
@@
LLaMA3-8B (Dubey et al., 2024) 100 81.0 72.5 31.5 59.0 61.6

### v3

@@ Abstract
[…]
@@ 1 Introduction
Since attention is introduced and becomes a fundamental component of Transformers (vaswani2017attention), multi-head attention has been the standard architecture for natural language processing (kenton2019bert) and computer vision tasks (dosovitskiy2021an). It is well known that using multiple heads can improve model accuracy. However, not all attention heads hold equal significance. Some works have shown that many attention heads can be pruned without affecting accuracy. For example, voita2019analyzing introduces a method to quantify the usefulness of each attention head and prune those that are redundant. Similarly, michel2019sixteen challenges the necessity of multiple heads by examining the impact of extensive pruning across various settings. In computer vision, some works also identify attention head redundancy. bhattacharyya2023decatt reduces redundancy to boost performance, while yun2024shvit develop single-head attention for efficiency. These findings demonstrate that vanilla multi-head attention contains redundant attention heads.
@@
To this end, we introduce Mixture-of-Head attention (MoH), a new architecture that integrates multi-head attention with the Mixture-of-Experts (MoE) mechanism (jacobs1991adaptive). Specifically, we propose to treat attention heads as experts within the MoE framework. Similar to MoE, MoH consists of multiple attention heads and a router that activates the Top-K heads for each token. Moreover, we replace the standard summation in multi-head attention with a weighted summation. This design offers two significant advantages: First, MoH allows each token to select the most relevant attention heads, improving inference efficiency without sacrificing accuracy or increasing the parameters. Second, by replacing the standard summation in multi-head attention with a weighted summation, MoH enhances the flexibility of the attention mechanism and increases the performance potential. Moreover, to efficiently capture common knowledge across different contexts, we designate a subset of attention heads as shared heads that remain always activated.
[…]
Furthermore, we demonstrate that pre-trained multi-head attention models, such as LLaMA3-8B (dubey2024llama), can be further continue-tuned into our MoH models. Specifically, using only about 3% (400B tokens) of the original LLaMA3 pre-training data for continue-tuning, MoH-LLaMA3-8B achieves an average accuracy of 64.0% across 14 benchmarks, outperforming LLaMA3-8B by 2.4% by utilizing only 75% of the attention heads. These results show that MoH is a promising alternative to vanilla multi-head attention, laying a solid foundation for developing advanced and efficient attention-based models. The main contributions are summarized as follows:
@@ 2 Related Work
Multi-Head Attention.   Transformers (vaswani2017attention) have garnered significant interest and success in both natural language processing and computer vision. The success of transformers has been long attributed to the multi-head attention mechanism (cordonnier2020multi). Multi-head attention mechanism is proposed by vaswani2017attention to enhance the representation power of an attention layer by allowing multiple attention heads to operate on different low-dimensional projections of the input. The outputs from these heads are then concatenated to form the final result. Alternatively, by decomposing the output projection matrix by rows, multi-head attention can be expressed in a summation form. In summation form, each head operates in parallel, and the final output is the sum of all heads. Inspired by this observation, we propose MoH, a dynamic attention-head routing mechanism that allows each token to adaptively select the appropriate heads.
@@
Mixture-of-Experts Models.   The Mixture-of-Experts (MoE) method (du2022glam; lewis2021base; rajbhandari2022deepspeed; roller2021hash; zhou2022mixture; jin2024moe++) is introduced to expand the capacity of deep neural networks without increasing computational costs. In this approach, only a subset of parameters, known as experts, is activated for each input. shazeer2017outrageously first introduces an MoE layer between LSTM layers. Switch Transformer (fedus2022switch) further simplifies the gating mechanism by selecting only the Top-1 expert per token. Gshard (lepikhin2020gshard) improves the Top-2 expert routing strategy. In contrast to MoE, which emphasizes efficient parameter scaling while maintaining manageable computational costs, our MoH focuses on reducing the activation of redundant attention heads without increasing the number of parameters.
@@
Attention Head Specialization and Efficiency.   Many recent studies show that not all attention heads in Transformers are equally useful. peng2020mixture proposes a mixture-of-heads approach, where only a few selected heads are used, yet the model performs just as well or even better. csordas2024switchhead pushes this further with SwitchHead, an MoE-style method that activates only a small number of heads for each token, speeding up inference while keeping performance high. In long-context language models, this idea is even more clear. wu2024retrieval shows that a few special retrieval heads are mainly responsible for keeping facts consistent in long inputs. fu2024not finds that keeping only the most useful heads in the KV cache can save memory. xiao2024duoattention proposes DuoAttention, which combines different types of heads to make long-context inference more efficient without losing quality. Similar patterns appear in vision models. gandelsman2023interpreting shows that CLIP’s attention heads each focus on specific visual features, and this can be explained through related text prompts. balasubramanian2024decomposing finds that this kind of head specialization also exists in other vision models beyond CLIP. basile2024residual shows that using only a few selected heads chosen by spectral methods can even beat the full model on zero-shot tasks.
@@ 3.1 Multi-Head Attention
We begin by reviewing the multi-head attention mechanism introduced by vaswani2017attention. The multi-head attention mechanism is based on scaled dot-product attention. Specifically, for $T$ tokens $\bm{X}\in\mathbb{R}^{T\times d_{in}}$ of $d_{in}$ dimensions each and $T^{\prime}$ tokens $\bm{X}^{\prime}\in\mathbb{R}^{T^{\prime}\times d_{in}}$ of $d_{in}$ dimensions each, the scaled dot-product attention is computed as follows:
@@
Concatenation Form.   To enhance the representation power, vaswani2017attention proposes to allow multiple attention heads to operate on different low-dimensional projections of the input tokens. Specifically, the multi-head attention mechanism computes $h$ different low-dimensional projections of $(\bm{Q},\bm{K},\bm{V})$, performs scaled dot-product attention for each head, concatenates the results, and applies a projection to the concatenated output. The concatenation form of the multi-head attention can be formulated as:
$\textrm{MultiHead}($ $\bm{X},\bm{X}^{\prime})=\textrm{Concat}(\bm{H}^{1},\bm{H}^{2},.,\bm{H}^{h})\bm{W}_{O},$ (2)
@@
Summation Form.   The multi-head attention mechanism is typically represented in its concatenation form. However, from another perspective, if we decompose $\bm{W}_{O}\in\mathbb{R}^{d_{v}\times d_{out}}$ by rows, we can express multi-head attention in a summation form. Specifically, $\bm{W}_{O}$ can be divided into $h$ matrices by rows, i.e., $\bm{W}_{O}^{1},\bm{W}_{O}^{2},...,\bm{W}_{O}^{h}]=\bm{W}_{O}$, where $\bm{W}_{O}^{i}\in\mathbb{R}^{d_{v}/h\times d_{out}}$. Finally, the summation form of the multi-head attention can then be formulated as:
@@
Table 1: Comparisons to current state-of-the-art methods on ImageNet-1K classification. Our MoH-ViT models, based on TransNeXt (shi2024transnext), are trained for 300 epochs using a resolution of $\times224$. To ensure a fair comparison, we only replace the standard multi-head attention with our Mixture-of-Head attention (MoH), keeping all other training parameters identical to TransNeXt.
@@
DeiT-S (touvron2021training) 22 100 79.8
T2T-ViT-19 (yuan2021tokens) 39 100 81.9
Swin-S (liu2021swin) 50 100 83.1
PVTv2-B3 (wang2022pvt) 45 100 83.2
CoAtNet-1 (dai2021coatnet) 42 100 83.3
Focal-S (yang2021focal) 51 100 83.5
FocalNet-S (yang2022focal) 50 100 83.5
MViTv2-S (li2022mvitv2) 35 100 83.6
UniFormer-B (li2023uniformer) 50 100 83.9
CAFormer-S36 (yu2023metaformer) 39 100 84.5
TransNeXt-S (shi2024transnext) 50 100 84.7
@@
DeiT-B (touvron2021training) 86 100 81.8
T2T-ViT-24 (yuan2021tokens) 64 100 82.3
Swin-B (liu2021swin) 88 100 83.5
PVTv2-B5 (wang2022pvt) 82 100 83.8
Focal-B (yang2021focal) 90 100 83.8
FocalNet-B (yang2022focal) 89 100 83.9
CoAtNet-2 (dai2021coatnet) 75 100 84.1
MViTv2-B (li2022mvitv2) 52 100 84.4
MOAT-2 (yang2022moat) 73 100 84.7
iFormer-L (si2022inception) 87 100 84.8
TransNeXt-B (shi2024transnext) 90 100 84.8
@@ 3.2 Mixture-of-Head Attention
Recently, the Mixture-of-Experts (MoE) method has emerged as a popular approach for scaling the parameters of large language models (jiang2024mixtral; muennighoff2024olmoe). A MoE layer consists of multiple expert networks and a router that activates the Top-K experts. Generally, the number of activated experts $K$ is significantly smaller than the total number of experts to ensure inference efficiency.
@@
Shared Heads.   In attention mechanism, some attention heads may capture common knowledge across different contexts, such as grammatical rules in language. Inspired by dai2024deepseekmoe, we designate a subset of heads as shared heads that remain always activated. By consolidating common knowledge within shared heads, we reduce redundancy among the other dynamically routed heads.
@@
$\alpha_{1},\alpha_{2}]=\text{Softmax}(\bm{W}_{h}\bm{x}_{t}),$ (6)
@@
Load Balance Loss   Directly training an MoE layer often causes the majority of tokens to be routed to a small number of experts, leaving the remaining experts insufficiently trained (shazeer2017outrageously). To avoid the unbalanced load in the proposed MoH, following previous MoE methods (lepikhin2020gshard; wei2024skywork), we apply a load balance loss. Specifically, for the $t_{th}$ input token $\bm{x}_{t}\in\mathbb{R}^{d_{in}}$ in $\bm{X}\in\mathbb{R}^{T\times d_{in}}$, the load balance loss $\mathcal{L}_{b}$ is formulated as:
@@ 4.1 ViT for Image Classification
Model Settings.   For Vision Transformers (ViT) (dosovitskiy2021an), our MoH-ViT models are implemented based on the TransNeXt (shi2024transnext) framework and trained from scratch on the ImageNet-1K dataset (deng2009imagenet), which contains over 1.2 million images in 1,000 categories. To ensure a fair comparison, we only replace the standard multi-head attention with the proposed MoH, while keeping all other training parameters identical to TransNeXt.
Training Details.   Our MoH-ViT models are trained for 300 epochs using automatic mixed precision across 8 GPUs. We follow the training strategy of TransNeXt, which includes various data augmentation techniques, including Random Augmentation (cubuk2020randaugment), Mixup (zhang2017mixup), CutMix (yun2019cutmix), and Random Erasing (zhong2020random). We also apply Label Smoothing (szegedy2016rethinking) and DropPath (huang2016deep) to regularize our models. We optimize our models using AdamW optimizer (loshchilov2017decoupled) with a gradient clipping norm of 1.0 and a weight decay of 0.05. The initial learning rate is set to 1e-3, with a 5-epoch warm-up starting at 1e-6. A cosine learning rate scheduler (loshchilov2016sgdr) is employed to decay the learning rate. During training, images are randomly cropped to a size of $\times224$. It is worth noting that we do not use Exponential Moving Average (EMA) weights.
Table 3: Comparisons between MoH-LLMs and vanilla LLMs.100B” denotes a training budget of 100 billion tokens, while “200B” denotes a budget of 200 billion tokens. We observe that larger models, e.g., MoH-LLM-B, generally perform worse than smaller models, e.g., MoH-LLM-S, on TruthfulQA, consistent with the findings reported by lin-etal-2022-truthfulqa.
@@
Methods #Activated Heads (%) HellaSwag (10) LogiQA BoolQ (32) LAMBADA SciQ
LLaMA3-8B (dubey2024llama) 100 81.9 30.0 83.9 75.5 94.0
@@
Methods #Activated Heads (%) PIQA WinoGrande NQ (32) ARC-C (25) Average
LLaMA3-8B (dubey2024llama) 100 81.0 72.5 31.5 59.0 61.6